OpenStudy (anonymous):

There are 20 passengers on the plane. Assume that any gender combination of the passengers is equally likely. What is the probability that there will be at least on woman on the plane?

OpenStudy (anonymous):

why is that?

OpenStudy (anonymous):

flip a coin 20 times, what is the likelihood that you will flip at least one heads? much more than 1/20

OpenStudy (anonymous):

im so confused!

OpenStudy (anonymous):

Going with the coin flip analogy, there is only one outcome that has all tails. all the rest have at least one heads. so, how many combinations were there?

OpenStudy (anonymous):

oh now i get it..goodness when i mess things like this up ik its time for bed...

OpenStudy (anonymous):

I get a ridiculous answer, so I am needing a check: 1048575/1048576

OpenStudy (zarkon):

that is correct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

which is (2^20-1)/(2^20) = 1048575/1048576

OpenStudy (zarkon):

P(at lest on F)=1-P(all M) =1-P(m1 and M2 and ... and M20) =1-P(m1)*P(m2)*...*P(M20) \[=1-\frac{1}{2}\times\frac{1}{2}\times\cdots\times\frac{1}{2}=1-\frac{1}{2^{20}}=\frac{2^{20}-1}{2^{20}}\] as already stated.

OpenStudy (zarkon):

*at least